OpenStudy (1645323):

Factor by grouping and solve:x3y2 - 8x3y + 16x3 + y5 - 8y4 + 16y3

OpenStudy (1645323):

@hartnn

OpenStudy (1645323):

@phi

OpenStudy (phi):

try the obvious (as a first step) it looks like you can factor x^3 out of the first 3 terms

OpenStudy (1645323):

so would it be x^3(y^2-8y+16+y^5-8y^4+16y^3)

OpenStudy (phi):

close, you put the parens around the terms where you *took out* the x^3 you don't put the parens around everything.

OpenStudy (phi):

try again. You almost have it. But you put the parens in the wrong places

OpenStudy (1645323):

(x^3)(y^2-8y+16+y^5-8y^4+16Y^3)

OpenStudy (phi):

I guess you are missing an idea. Here is another way to think about it. if you have a( a +b) +c and you distribute the "a" (multiply each term inside the parens by a) you get aa +ab + c or a^2 +ab + c if you "undo that" (factor out the a) you take an "a" out of a^2 and out of ab leaving a+b we put the a out front: a(a+b) + c notice we don't put the parens around the c a(a+b+c) means aa +ab+ac and that is different from aa+ab+c

OpenStudy (1645323):

oh so x^3(y^3)-(8y)+(16)+y^5-8y^4+16y^3

OpenStudy (phi):

there are 3 terms where you "took out" x^3 you put all 3 terms in one set of parens. Think like this: there is a package of terms where I took out the x^3. Put those terms in one "bag" and label the bag with x^3 on the outside to show we took out the x^3

OpenStudy (phi):

in other words, you use one ( and one )

OpenStudy (phi):

that you use to show the 3 terms where you took out the x^3

OpenStudy (phi):

not clear ?

OpenStudy (1645323):

ok im so dumb x^3(y^2-8y+16)+y^5-8y^4+16y^3

OpenStudy (phi):

exactly. As a way to check the answer, can you distribute the x^3 ? what do you get ?

OpenStudy (1645323):

just like th original equation

OpenStudy (phi):

Yes. The main idea is to see that \[ x^3(y^2-8y+16) \] and \[ x^3y^2 - 8x^3y + 16x^3 \] mean the something, and you can switch between them (factor or distribute) that idea has to be clear in your mind.

OpenStudy (phi):

*same thing (not something )

OpenStudy (phi):

now back to the problem \[ x^3(y^2-8y+16)+y^5-8y^4+16y^3 \] do you see what you can factor out the last 3 terms ?

OpenStudy (1645323):

y^3

OpenStudy (phi):

yes, and if you factor y^3 out of the last 3 terms, what does the whole expression become ?

OpenStudy (1645323):

x^3(y^2-8y+16) +y^3(y^2-8y+16)

OpenStudy (phi):

looking good. now, though (y^2-8y+16) looks complicated, it is one "thing" and can be factored out of each term

OpenStudy (1645323):

2

OpenStudy (1645323):

4

OpenStudy (phi):

we don't have to factor (y^2-8y+16) yet. Instead, if you had this problem \[ x^3 a + y^3 a \] could you factor it ? what do you get ?

OpenStudy (1645323):

x^3 4 +y^3 4

OpenStudy (phi):

I don't know how you are getting a 4. We are working on your problem, but to make progress I want to do a simpler problem first the simpler problem is \[ x^3 a + y^3 a \] what is in both terms that you can factor out ?

OpenStudy (1645323):

3

OpenStudy (phi):

the only 3 we have in \[ x^3 a + y^3 a \] are the exponents (the little 3 in the upper right) we don't factor out exponents... they are just a short-cut to write \[ x\cdot x \cdot x \cdot a + y\cdot y \cdot y \cdot a \]

OpenStudy (phi):

there is only one variable (letter) that you can factor out of each term in \[ x\cdot x \cdot x \cdot a + y\cdot y \cdot y \cdot a \]

OpenStudy (1645323):

x and y

OpenStudy (phi):

x is *only in the first term* y is *only in the second term* you need to find a variable that is in *both terms* that you can "take out"

OpenStudy (1645323):

a

OpenStudy (phi):

yes, only "a" can be factored out. what do you get ?

OpenStudy (1645323):

x^3y^3(a)

OpenStudy (phi):

almost. you start with \[ x^3 a + y^3 a \] you keep the plus sign (you start with two terms and you still have two terms (in parens) when you are done.

OpenStudy (1645323):

x^3+y^3(a)

OpenStudy (1645323):

@phi

OpenStudy (phi):

ok, but where does the ( ) go ? remember, the ( ) go around the terms where you took out the a. The ( ) show which terms used to have the a.

OpenStudy (1645323):

(x^3+Y^3)a

OpenStudy (phi):

yes. Does that make sense ?

OpenStudy (1645323):

ys thansk

OpenStudy (phi):

The reason we went over that is that this rule works even if "a" is complicated for example, if a was (y^2-8y+16) it works the same way. can you factor \[ x^3 (y^2-8y+16) + y^3 (y^2-8y+16) \]

OpenStudy (phi):

do you see you can take out (y^2-8y+16) from both terms ?

OpenStudy (1645323):

yes @phi

OpenStudy (phi):

what do you get ?

OpenStudy (1645323):

X^3+y^3

OpenStudy (phi):

that is what you get after you take out (y^2-8y+16) but you have to put that in parens. and put (y^2-8y+16) on the outside of the parens (otherwise you changed the problem)

OpenStudy (phi):

in other words, your problem can be factored into \[ ( x^3 + y^3) (y^2-8y+16) \]

OpenStudy (phi):

y^2 -8y+16 factors into (y-4)^2 so the final answer is \[ (x^3 + y^3) (y-4)^2 \]